1. Graph 2 20 16+ Marginal cost 127 Price ($) ATC MR Demand = WTP 100 200 20 - Quantity (Q)
  2. a- (5 pts) What is the maximum possible profit that the monopolist can earn? (Calculate the exact amount of profit).
  1. (5 pts) How many units of output should a monopolist produce if the monopolist wants to maximize revenue rather than profits?

  1. (5 pts) What is the maximum obtainable total revenue for the firm?

  1. (5 pts) At the level of output you identified in part g, how much profit would the firm earn?
Answer #1

a)

A monopolist maximizes profit by producing such that Marginal Revenue=Marginal Cost.

We can see that Marginal Cost and Marginal Revenue Curves intersect each other at

Price=$12

Quantity=80

ATC at Q=80 is $2 (Refer given graph)

Maximum Profit=(P-ATC)*Q=(12-2)*80=$800

b)

At maximum Revenue, MR should be zero. MR cuts horizontal axis at Q=100

Monopolist should produce 100 units to maximize Revenue.

c)

Refer to above demand curve, Price,P =$10 for Q=100

Maximum Revenue=P*Q=10*100=$1000

d)

At output level of 100, ATC=$2.5 (Approx)

Profit at this level=(P-ATC)*Q=(10-2.5)*100=$750